OpenStudy (anonymous):

For which system of equations is (–2, 2) the solution?

OpenStudy (darkigloo):

are there any answer choices?

OpenStudy (anonymous):

. 3x – 5y = –16 –2x + 3y = 10 B. –4x + 2y = 6 2x – 3y = 10 C. 4x + 3y = 2 5x + 4y = –2 D. 6x + 5y = –2 –8x + 5y = –6

OpenStudy (anonymous):

plug in the numbers to the equation